at the radius of the earth's orbit, the intensity of the sunlight is I=1.4kW/m².(this means that 1.4 kW/m² passes thorough 1 square metre at right angles to the sun's rays.) the solar array consist of 10 panels, each with area A= 1.6m². under optional conditions -with the sun at the right angles to the array and no clouds in the sky-the array produces Power output =2.3kW. what is the efficiency of the array? (power output/power input-------%)
1
Expert's answer
2020-05-18T10:21:54-0400

Let us calculate the total amount of energy that passes through the array per 1 second:

"I\cdot A\cdot 10 = 1.4\,\mathrm{kW/m^2}\cdot1.6\,\mathrm{m}^2\cdot10 = 22.4\,\mathrm{kW}."

If the total output is 2.3 kW, we can calculate the efficiency

"\eta = \dfrac{2.3\,\mathrm{kW}}{22.4\,\mathrm{kW}} \approx 0.10 = 10\%."